For the following reaction, 5.80 grams of sulfuric acid are mixed with excess calcium hydroxide. The reaction yields 5.77 grams of calcium sulfate.
1
Expert's answer
2020-01-23T06:02:10-0500

H2SO4+Ca(OH)2"\to" CaSO4+2H2O

n(H2SO4)=5.8(g)/98.078(g/mol)=0.059(mol)

m(CaSO4)=136.14(g/mol)*0.059(mol)=8.032(g)

experimental yield =5.77(g)/8.032(g)*100%=71.84%